OpenStudy (anonymous):

Assume that Theta is an acute angle in a right triangle satisfying the given conditions. Evaluate the indicated trig function.

OpenStudy (anonymous):

\[\csc \theta = \frac{ 15 }{ 2 }; \cot \Theta\]

OpenStudy (anonymous):

if you want the angle you would use soh cah toa. you know that the hypotenuse is going to be across from the 90 degree side since on a 90 degree triangle it is always going to be the largest angle possible for that triangle.

OpenStudy (anonymous):

 So if csc is \[\frac{ 1 }{ \sin x }\] that would make it \[\frac{ 1 }{ 15/2 }\] no? Maybe I'm just thinking too much.

OpenStudy (campbell_st):

define the angle and then use the information to indentify the sides  find tan(x) then take the reciprocal for cot(x)